OpenStudy (anonymous):

Find functions f,g:N-->N, so that the following are true at the same time: 1. f is not surjective 2. g is not injective 3. g o f is bijective I have been trying to figure this out for hours and nothing seems to work!!

geerky42 (geerky42):

Not sure, but how about these functions? \(f(x) = 2+\sqrt {x-1}\\~\\g(x) = (x-2)^2+1\)

geerky42 (geerky42):

Maybe somehow have function be piecewise-defined?

OpenStudy (anonymous):

g(x) isn't surjective either though (eg you can't get g(x)=3).. so how is gof gonna be bijective?

geerky42 (geerky42):

Well question never said that \(g\) has to be surjective. Also \((g\circ f)(x) = x\) for \(x\ge1\), which makes it bijective, right?

OpenStudy (anonymous):

well g(x) necessarily has to be surjective for gof to be surjective....

OpenStudy (anonymous):

I mean definition says that for gof to be bijective f has to be injective and g surjective

geerky42 (geerky42):

Well I'm out of luck. Maybe these users can help: @amistre64 @TuringTest @perl Or you can ask this question on math stack exchange?

OpenStudy (anonymous):

Will give that a go, thanks!

OpenStudy (amistre64):

bijective means 1-1 and onto surjective in onto, injective is 1-1 just trying to remember the defs

OpenStudy (amistre64):

f is not surjective; or f is injective (1-1 but not onto) f(n) = 2n is clearly 1-1 but not onto

OpenStudy (amistre64):

surjective is onto, but need not be 1-1 simplest thing is to right shift a parabola g(n) = (n-1)^2 , does n include 0? if not up it by 1

OpenStudy (amistre64):

oh, you are trying to find f and g such that all 3 conditions hold at the same time

OpenStudy (xapproachesinfinity):

the question was to find g(f(x)) which is one to one correspondence regardless of f and g i guess

OpenStudy (amistre64):

g(f) = n + k is superb for bijective lol

OpenStudy (amistre64):

well, k=0 maybe

OpenStudy (amistre64):

g and f should be inverses then, or at least one way inverses